Bond Prices Question 6.

Mutt Inc has 5.5% coupon bonds that have 3.5 years to maturity and a price of $978.56. What is the yield to maturity?

Mutt has another bond with a 4.75% coupon with 6 years to maturity and a yield to maturity of 3.94%. What is the price of the bond?

GW Corp issued a 10-year bond 2 years ago with a coupon rate of 5.8%. If the yield to maturity on the bond is 6.6%, what is the current price?

GW Corp also has just issued an 8-year discount bond for an issue price of $661.56. What is the yield to maturity?
Answer #1

6. a) Coupon =5.5%*1000 =55
Number of years =3.5
PV =-978.56
YTM using excel formula =RATE(3.5,55,-978.56,1000) =6.20%

b) Coupon =4.75%*1000 =47.50
Number of Periods =6
YTM =3.94%
Price of Bond =PV of Coupons+PV of Par Value =47.50*((1-(1+3.94%)^-6)/3.94%)+1000/(1+3.94%)^6 =1042.54

c)Number of Periods =10-2 =8
Coupon =5.8%*1000 =58
YTM =6.6%
Price of Bond =PV of Coupons+PV of Par Value =58*((1-(1+6.6%)^-8)/6.6%)+1000/(1+6.6%)^8 =951.48

d) YTM =(1000/661.56)^(1/8)-1 =5.30%